For the graph of the function f (x) = x2 – 5x + 6, find the vertex and the axis of symmetry. I got vertex=(5/2,-1/4) x= 5/2

Found 2 solutions by ewatrrr, nerdybill:
Answer by ewatrrr(24785)

 
Hi
f(x) = x^2 – 5x + 6 = (x-5/2)^2 - 25/4 + 24/4 = (x-5/2)^2 - 1/4
Yes! V(5/2,-1/4) and x = 5/2 the line of symmetry
x - intercepts: 0 = (x-5/2)^2 - 1/4 x-5/2 = ± 1/2
(3,0) & (2,0)


Answer by nerdybill(7384)

You would be correct.